OpenStudy (anonymous):

Solve 1 = cot^2(x) = 8sinx, for 0(lesser-than-or-equal-to) x (lesser-than-or-equal-to) 360 degrees??

OpenStudy (anonymous):

r u sure u have correct question? The two '=' signs?

OpenStudy (anonymous):

sorry, 1+ cot^2.....

OpenStudy (anonymous):

my bad. didn't press shift.

OpenStudy (anonymous):

ahh u could try replacing cot^2x with cosec^2x - 1 = 1/sin^2x + 1

OpenStudy (anonymous):

but that would be 1 + coesc^2x - 1, which would be cosec^x.

OpenStudy (anonymous):

hmm lets try this: 1 + ( 1 - sin^2x) = 8 sinx ---------- sin^2x sin^2x + 1 - sin^2x = 8 sin^3 x sin^3 x = 1/8 s in x = +/- 1/2

OpenStudy (anonymous):

x = 30, 150, 210, 330 degrees

OpenStudy (anonymous):

how did you get that?? \[1 + \cot^2x = 8sinx\] so \[cosec^2x = 8sinx\] now what??

OpenStudy (anonymous):

cosec^2 x = 1 / sin^2 x

OpenStudy (anonymous):

yes. so?

OpenStudy (anonymous):

you replace cot^2x by cosec^2 x - 1 because cosec^2x = 1 + cot^ 2 x

OpenStudy (anonymous):

but you can replace 1 = cot^2x by cosec^2x straight away!!

OpenStudy (anonymous):

*1 + cot^2x

OpenStudy (anonymous):

yea right - i didnt see the obvious! - but my method comes out with the same answer. your way is better though! they both come to 1 / sin^2x = 8 sinx cross multiply: 8 sin^3 x = 1 sin^3 x = 1/8 sin x = 1/2 or -1/2

OpenStudy (anonymous):

ohhh. cool. i didn't know how to simplify it :P

OpenStudy (anonymous):

but it can't be -0.5, though, can it??

OpenStudy (anonymous):

hold on - sorry - i-m prone to error today the cube root of 1/8 is 1/2 only so x = 30, 150 only

OpenStudy (anonymous):

i think tahs correct now lol!

OpenStudy (anonymous):

it's okayy, i have these weird days too :P

OpenStudy (anonymous):

can you help me with another question??

OpenStudy (anonymous):

ok